Booker, Inc., has identified an investment project with the following cash flows. Year Cash Flow $1,000 1,230 1,450 2,190 m IA step by step explanation would be greatly appreciated! I am not the best with excel or a calculator so a walk through to the answers would be helpful!
Answer #1

FV = PV(1 + r)^t FV @ 9% = $1,000(1.09)^3 + $1,230(1.09)^2 + $1,450(1.09) + $2,190 = $6,526.89 FV @ 12% = $1,000(1.12)^3 + $1